OpenStudy (anonymous):

How to find y from differential equation: (y'')^2+y'=xy'' ?:)

OpenStudy (anonymous):

maby, substitution y'=yz and y''=y(z'+z^2)?

OpenStudy (anonymous):

notice there are no y term u= y' u' ^2 + u=x u' u' + u/u'=x

OpenStudy (anonymous):

but then..?

OpenStudy (anonymous):

...u=u'(x-u').. have any idea? :)

OpenStudy (anonymous):

I have no idea, but here is how wolfram solves it http://www4a.wolframalpha.com/Calculate/MSP/MSP20719ibiid3766b33e500000g3d326gh6486h02?MSPStoreType=image/gif&s=38&w=350&h=863

OpenStudy (anonymous):

tnx for it ;)

OpenStudy (jamesj):

Nice problem.

OpenStudy (anonymous):

\[(y'')^2+y'=xy'' \] is a nonlinear 2nd order differential equation. Some type of substitution is needed to put it in a linear form.

OpenStudy (jamesj):

@rolands, you have the answer now, yes?

OpenStudy (anonymous):

no..

OpenStudy (jamesj):

Ok, so here it is. Using the substitution suggested where u = y', we have \[ (u')^2 + u = xu' \] Differentiate now that equation and we have \[ 2 u' u'' + u' = u' + xu'' \] i.e., \[ u'' (2u' - x) = 0 \] Consider now each case in turn...

OpenStudy (jamesj):

If u'' = 0, then u = C1.x + C2 If 2u' = x, then u = x^2/4 + C3 Now integrate u once more to recover y.

OpenStudy (anonymous):

tnx ;)